OpenStudy (anonymous):

(13,20)and(18,8) Find the distance between the points

OpenStudy (anonymous):

Distance between points (x1,y1) and (x2,y2) is

OpenStudy (anonymous):

difference of y-co-ordinates = 20-8 = 12 its square will be 144

OpenStudy (uri):

Distance formula:\[\sqrt{(x2-x1)^2 +(y2-y1)^2}\]

OpenStudy (anonymous):

according to the formula, you need to square it , so (-5)^2 =... ?

OpenStudy (uri):